OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)) y=4/(x^2)+9

OpenStudy (anonymous):

\[y=\frac{ 4 }{ x ^{2}}+9\]

OpenStudy (anonymous):

ok that changes it

OpenStudy (anonymous):

you could say \(f(x)=x+9\) and \(g(x)=\frac{4}{x^2}\) but there are other ways too

OpenStudy (anonymous):

oh ok:) i got it thanks :)